# RDA-INR: Riemannian Diffeomorphic Autoencoding via Implicit Neural Representations

### v2

@@ 4.4 Encoding objects
$\min_{z}\mathcal{D}_{rec}((\phi_{1}^{z})^{-1}\cdot\mathcal{T}_{\theta},\mathcal{S})+\gamma\lVert z\rVert_{2}^{2},$ (18)
where $\gamma\in\mathbb{R}$, $\phi_{t}^{z}$ is defined via the ODE in optimization problems (14) and (15), $\mathcal{T}_{\theta}$ is the learned template, and $\mathcal{D}_{rec}$ is some reconstruction data-fitting term. For instance, in case $\mathcal{S}$ is represented by a mesh or a point cloud, we use $\operatorname*{\mathop{\mathbb{E}}}_{x\in\mathcal{S}}\left[\lvert I(x,z,1)\rvert\right]$ as $\mathcal{D}_{rec}$ with $I(x,z,t)$ given by Equation (13).
Remark 2.
@@
$\pi_{H}(p)=\operatorname*{arg\,min}_{q\in H}d(p,q)^{2},$
where $d$ is the Riemannian distance on the Riemannian manifold $M$. Alternatively, this can be reformulated into a minimization over the subspace $W$:
@@
Equation (18) resembles this optimization problem. First, we have $p=\mathcal{S}$. Furthermore, $(\phi_{1}^{z})^{-1}\cdot\mathcal{T}_{\theta}$ should approximate $\text{exp}_{\mu}$ in case we consider the LDDMM manifold with $\mu=\mathcal{T}_{\theta}$. Although $\mathcal{D}_{rec}$ is not a Riemannian distance, it replaces the Riemannian distance $d$, as also done in, e.g., Charlier et al. [47]. Finally, instead of searching over the vector space $W$, we search over a latent space that defines a point $(\phi_{1}^{z})^{-1}\cdot\mathcal{T}_{\theta}$ on the manifold. Hence, instead of finding an initial velocity $w\in W$ to define a point $q=\text{exp}_{\mu}(w)$ on the manifold, we find a latent code $z$ that determines a point on the manifold.
@@ 5 Numerical results
In this section, we demonstrate the benefit of the physical consistency and resolution independence of our model. To evaluate the benefit of physical consistency, we compare the Riemannian LDDMM regularizer to a non-Riemannian regularizer. We utilize two shape datasets for this comparison: a synthetic rectangles dataset and a liver dataset [21, 86]. Subsequently, we compare our method to DAE [16] on the liver dataset to illustrate the benefit of resolution independence for LDDMM statistical latent modeling.
@@ 5.1 Physical consistency and statistical shape modeling via latent spaces
To assess whether physical consistency is desirable for statistical shape modeling via latent spaces, we compare our model that regularizes the flow via the Riemannian LDDMM regularization (see Equation (15)) to our model utilizing the non-Riemannian pointwise regularization, which is presented in Appendix D.2. The latter model is the state-of-the-art baseline model by Sun et al. [21] with a different data-fitting term and a different approach to solving the ordinary differential equation.
We assess the mean-variance analysis capabilities of the models by visualizing the learned templates and deformations from the template to the training shapes. We relate our findings to Figure 2. Subsequently, we discuss reconstruction generalization and robustness of the reconstruction procedure to noisy data. This discussion shows the effect of Riemannian regularization on the quality and stability of the reconstruction procedure. We also relate this discussion to Figure 2. For details about the data source, the data preprocessing, and the training, we refer to Appendix D.
@@ 5.1.1 Learning shape Fréchet means
(a) Liver Riemannian regularization ($\eta=50$).
@@
(c) Rectangles Riemannian regularization ($\eta=0.05$).
(d) Rectangles non-Riemannian regularization.
@@
Figure 5: The learned templates of two different models trained on the rectangles dataset and the liver dataset. We use the model learned using the Riemannian LDDMM regularization ((a) and (c)) and the model learned using the non-Riemannian pointwise loss ((b) and (d)). We use $\eta=0.05$ and $\eta=50$ in Equation (16) for the rectangles and liver dataset, respectively.
@@
For training the models with Riemannian LDDMM regularization, we need to pick a value for $\eta$. We choose $\eta=0.05$ for the rectangles dataset as we expect rigid body motions from the template to the training shapes. For the liver data, the only prior that we have is that the liver should not change too quickly. Hence, we pick $\eta=50$ such that $\lVert\cdot\rVert_{V}\approx\eta\lVert\cdot\rVert_{L_{2}(\Omega)}$.
@@
After training, we obtain the template shapes in Figure 5. The templates obtained via the Riemannian regularization and the non-Riemannian regularization look almost identical for the liver dataset. The reason is that there is no clear deformation prior for the liver dataset besides that the liver should not change too quickly. The non-Riemannian regularization induces this prior by construction while the Riemannian regularization enforces this by our choice of $\eta=50$. More precisely, as $\lVert\cdot\rVert_{V}\approx\eta\lVert\cdot\rVert_{L_{2}(\Omega)}$ and as the pointwise loss can be interpreted as a non-Riemannian version of the LDDMM regularization with $\lVert\cdot\rVert_{V}=\lVert\cdot\rVert_{L_{2}(\Omega)}$ (see Appendix D.2), both regularizations induce the same prior on the deformation. Hence, we expect similar templates in Figures 4(a) and 4(b).
@@
While the templates of the liver dataset look the same, the templates for the rectangles dataset differ. This difference between the rectangles and the liver dataset stems from the difference in the prior. While the liver shapes should only change gradually, the rectangles should also move rigidly. Consequently, we let the Killing energy play a more prominent role in $\lVert\cdot\rVert_{V}$ in the Riemannian LDDMM regularization. This choice yields almost rigid deformations from the template to the training shapes, which results in the square template in Figure 4(c). As the pointwise loss does not induce an isometry prior and is a non-Riemannian version of the LDDMM regularization with $\lVert\cdot\rVert_{V}=\lVert\cdot\rVert_{L_{2}(\Omega)}$, we obtain the spherical template in Figure 4(d). This template is not desired as it lies outside the data distribution, as depicted in the third category in Figure 2.
@@
In summary, while on the liver dataset, the templates look similar, the Riemannian regularization is required for the rectangles dataset to learn a template that fits the data distribution. Hence, the Riemannian LDDMM regularization allows for more flexibility in learning a template shape that resembles the data.
@@ 5.1.2 Learning geodesic shape deformations
Figure 6: The deformation of the template into a reconstructed rectangle using the model learned using the Riemannian LDDMM regularization (top) and using the model learned using the non-Riemannian pointwise loss (bottom). The colors represent the matching of the points to their template.
In the previous section, we showed the effect of a regularizer on the learned Fréchet mean. To finalize the discussion on Figure 2, we assess the variance calculation by qualitatively assessing whether the template deformations are geodesics.
In Figure 6 the template is transformed into a reconstructed training shape for each model trained on the rectangles dataset. We notice that the non-Riemannian model quickly transitions from the spherical template to the target shape and then has a tendency to stay in a nearly identical configuration. In contrast, the model with the Riemannian LDDMM regularization smoothly rotates and scales the rectangular template shape to the final reconstructed shape. Hence, it fits the rigid body motion prior induced by the Killing energy. As the non-Riemannian pointwise loss does not penalize big abrupt deformations and does not use a rigid body motion prior, we obtain non-smooth and non-rigid deformations in this case.
@@
Figure 7: The deformation of the template into a reconstructed liver using the model learned using the Riemannian LDDMM regularization (top) and using the model learned using the non-Riemannian pointwise loss (bottom). The colors represent the matching of the points to their template.
Figure 8: The velocity vector fields at $t=0$ and $t=0.2$ of the deformations in Figure 7. The colors represent the magnitude of the velocity vector field. At $t=0.2$, only the vector field at the tip of the shape is shown.
@@
Regarding the liver dataset, Figure 7 shows that the deformation with the Riemannian LDDMM regularization is smoother in time than the deformation with the non-Riemannian pointwise loss. Figure 8 reinforces this as it shows that the vector field of the non-Riemannian model at time $t=0.0$ is everywhere much larger than the vector field of the Riemannian model. Moreover, at $t=0.2$, the vector field of the Riemannian model points in approximately the same direction as at $t=0.0$, while this does not apply to the non-Riemannian model. Hence, the vector field of the Riemannian model is smoother in time.
@@
To explain these observations, we note that the non-Riemannian pointwise loss encourages a rapid transition to the target shape followed by a tendency to stay in a nearly identical configuration. Consequently, the velocity vector fields at later times $t$ fine-tune the obtained reconstructions. Hence, the velocity vector field at $t=0.2$ does not necessarily point in approximately the same direction as the velocity vector field at $t=0.0$, as shown in Figure 8. In contrast, with the Riemannian LDDMM regularization, rapid transitions are penalized, and it is preferred to gradually get closer to the target shape. Consequently, the Riemannian regularization allows for a smooth, physically plausible template deformation into the reconstructed shapes.
@@
In Table 2, we summarize the discussion regarding Fréchet means and variance. The table shows that the Riemannian LDDMM regularization is more flexible in influencing the template than the non-Riemannian pointwise loss. Moreover, to obtain good variance estimates, the LDDMM regularization is needed to obtain geodesic deformations of the template to the reconstructions. Hence, the model with the Riemannian LDDMM regularization is the only model in category 1 of Figure 2 and, therefore, is the model with the best mean-variance analysis. The Riemannian model allows us to calculate Fréchet means of the data, to calculate physically-plausible geodesic deformations between the template shape and another shape, and to approximate the Riemannian distance between the template and a target shape. Hence, we have added Riemannian geometry to the latent space model.
@@
Table 2: PGA performance. A summary of the results of Sections 5.1.1 and 5.1.2 regarding the models with Riemannian LDDMM regularization and non-Riemannian pointwise regularization. The results relate to Figure 2.
@@
Good Fréchet mean Good variance
@@
5.1.3 Generalizability of shape encoding
This section evaluates the reconstruction quality of the two different models by reconstructing the test sets of the rectangles and liver dataset. We use the Chamfer Distance (CD) and the Earth Mover Distance (EM) as evaluation metrics. Table 3 shows the reconstruction metrics of the training data. We see that both models perform approximately equally on the training data.
Table 3: Training set evaluation. The model with the Riemannian LDDMM regularization and the model with the non-Riemannian pointwise regularization are evaluated by reconstructing the training sets of the rectangles (Rect.) and liver dataset. We calculate the average value and median (between brackets) of the Chamfer Distance (CD) and Earth Mover distance (EM). The Chamfer Distance values are of the order $10^{-4}$. The best (smallest) values are in bold.
@@ 5.1.3 Generalizability of shape encoding
Table 4 shows the reconstruction metrics on the test sets. We can immediately see that on the rectangle dataset, the model with Riemannian LDDMM regularization performs the best. On the liver dataset, the method with Riemannian regularization and the method with non-Riemannian pointwise regularization perform approximately equally.
Table 4: Test set evaluation. The model with the Riemannian LDDMM regularization and the model with the non-Riemannian pointwise regularization are evaluated by reconstructing the test sets of the rectangles (Rect.) and liver dataset. We calculate the average value and median (between brackets) of the Chamfer Distance (CD) and Earth Mover distance (EM). The Chamfer Distance values are of the order $10^{-4}$. The best (smallest) values are in bold.
@@
The previous observations are related to the learned Fréchet mean. In the case of the rectangles dataset, we saw that the model with the Non-Riemannian pointwise loss did not learn a Fréchet mean resembling the data. As a consequence, the deformations of the spherical template do not end in proper rectangles. Hence, the fact that the model falls into category 3 in Figure 2 might cause generalization problems for the pointwise loss model on the rectangles dataset. This view is strengthened by the results on the liver dataset where the pointwise loss model performs the best. In that case, a physically plausible Fréchet mean is learned and the deformations end at a liver shape again. To summarize, a good Fréchet mean improves generalization performance. Since we have already shown that physical consistency improves the Fréchet mean, physical consistency can also enhance generalization performance.
@@ 5.1.4 Robustness to noise
In this section, we investigate how noise affects the reconstruction performance of the two models. We add random Gaussian noise with mean zero and a standard deviation of $0.01$ or $0.02$ to the vertices of the meshes in the test set. Subsequently, we reconstruct these noisy meshes and compare the reconstructions to the noiseless ground truth meshes. Figure 9 provides some of these reconstructions. In general, both models are effective in denoising the illustrated input meshes. On the rectangles, the Riemannian regularization produces good, noise-resistant reconstructions. The model with the non-Riemannian regularization generates worse reconstructions and is affected by the noise, as illustrated by the reconstructions, which gradually lose their rectangular shape when the noise increases. On the liver dataset, both models produce satisfactory denoised reconstructions. However, while the model with Riemannian regularization yields reconstructions that are approximately independent of the noise level, the model with non-Riemannian regularization results in visibly different reconstructions when the noise level varies.
Figure 9: The reconstruction of two test shapes with added zero-mean Gaussian noise for different standard deviations $\delta$. The red and blue colors represent the matching of the points to the learned template.
To assess whether the observations concerning Figure 9 extend to other liver and rectangle shapes, Table 5 presents the average and median reconstruction errors between the reconstructed shapes and the ground truth noiseless shapes. Comparing the values to Table 4, we notice that on the rectangles dataset, the model with Riemannian LDDMM regularization is much less sensitive to the noise than the model with the non-Riemannian regularization. On the liver dataset, the median values are relatively stable for both models. Moreover, the average reconstruction errors of both models are not influenced much by a noise level of $0.01$. However, when increasing the standard deviation to $0.02$, the non-Riemannian model attains a much bigger average reconstruction error. In contrast, the Riemannian model attains a similar average error to the scenario with a $0.01$ noise level.
Table 5: Noisy test set evaluation. The model with the Riemannian LDDMM regularization and the model with the non-Riemannian pointwise regularization are evaluated by reconstructing noisy test sets ($\delta=0.01$ and $\delta=0.02$) of the rectangles (Rect.) and liver dataset. We calculate the average value and median (between brackets) of the Chamfer Distance (CD) and Earth Mover distance (EM). The values are obtained by comparing the noiseless shapes to the reconstructions of their noisy versions. The Chamfer Distance values are of the order $10^{-4}$. The two best (smallest) values are in bold.
@@
The latter observation can be related to Figure 9 where the liver reconstructions obtained via the non-Riemannian model are dependent on the noise level. As discussed in Section 5.1.2, the non-Riemannian regularization encourages a rapid transition to the target shape followed by smaller deformations for fine-tuning. Such large displacements can cause sensitivity to small changes and might yield worse or different reconstructions. Disallowing such large displacements by applying the Riemannian regularization to the velocity vector fields stabilizes the problem.
Finally, we again notice a difference between the rectangle and liver dataset, which again stems from the learned Fréchet mean. In the case of the rectangles dataset, the pointwise loss model does not learn a physically plausible Fréchet mean of the data and the end of the template deformations do not correspond to test rectangles. In contrast, on the liver dataset a physically plausible template is learned and the end of the template deformations do correspond to livers. These observations strengthen our finding, as discussed in Section 5.1.3, that a physically plausible Fréchet mean, and hence LDDMM physical consistency, enhances generalization performance.

### v3

@@ 4.3 Encoding objects
(16) $\min_{z}\mathcal{D}_{rec}((\phi_{1}^{z})^{-1}\cdot\mathcal{T}_{\theta},\mathcal{S})+\gamma\left\lVert z\right\rVert_{2}^{2},$
where $\gamma\in\mathbb{R}$ corresponds to an isotropic Gaussian prior on the latent codes, $\phi_{t}^{z}$ is defined via the ODE in optimization problems (12) and (14), $\mathcal{T}_{\theta}$ is the learned template, and $\mathcal{D}_{rec}$ is $\operatorname*{\mathop{\mathbb{E}}}_{x\in\mathcal{S}}\left[\lvert I(x,z,1)\rvert\right]$ with $I(x,z,t)$ given by Equation (11) and $\mathcal{S}$ is represented by a mesh or a point cloud. Note that this $\mathcal{D}_{rec}$ differs from $\mathcal{D}$ in Equation (13). As the latent space generates a prior on the data, reconstructing solely with the point cloud vertices is sufficient.
Remark 4.3.
@@
Equation (16) resembles this optimization problem. Similarly to the relationship between Equation (12) and LDDMM PGA, $(\phi_{1}^{z})^{-1}\cdot\mathcal{T}_{\theta}$ should approximate the exponential map and instead of searching over a linear vector space $W$, we search over a nonlinear submanifold of vector fields defined via a latent space. Finally, although $\mathcal{D}_{rec}$ is not a Riemannian distance, it replaces the Riemannian distance $d$, as also done in, e.g., Charlier et al. [17].
@@ 5 Numerical results
In this section, we demonstrate the benefit of the physical consistency and resolution independence of our model. To evaluate the benefit of physical consistency, we compare the Riemannian LDDMM regularizer to a non-Riemannian regularizer. We utilize two shape datasets for this comparison: a synthetic rectangles dataset and a liver dataset [63, 21]. Subsequently, we compare our method to the diffeomorphic autoencoder (DAE) [14] on the liver dataset to illustrate the benefit of resolution independence for LDDMM statistical latent modeling. For details about the data source, the data preprocessing, and the training, we refer to Appendix D.
@@ 5.1 Physical consistency and statistical shape modeling via latent spaces
To assess whether physical consistency is desirable for statistical shape modeling via latent spaces, we compare our model that regularizes the flow via the Riemannian LDDMM regularization (see Equation (14)) to our model utilizing the non-Riemannian pointwise regularization, which is presented in Appendix D.2. The latter model is the state-of-the-art baseline model by Sun et al. [63] with a different data-fitting term and a different approach to solving the ODE. In Appendix B.2, we show that our changes do not deteriorate the reconstruction performance. Furthermore, in Appendix D.2, we explain the relationship of the non-Riemannian regularization with the Riemannian LDDMM regularization and explain why we call it a non-Riemannian regularization. Essentially, the pointwise regularization, like our Riemannian regularization, penalizes deviations of the time-dependent diffeomorphism at the final time from the identity diffeomorphism. However, the intermediate time diffeomorphisms are not promoted to correspond to a Riemannian geodesic between the identity and final diffeomorphism.
We assess the mean-variance analysis capabilities of the models by visualizing the learned templates and deformations from the template to the training shapes. We relate our findings to Figure 2. Subsequently, we discuss reconstruction generalization and robustness of the reconstruction procedure to noisy data. This discussion shows the effect of Riemannian regularization on the quality and stability of the reconstruction procedure.
@@ 5.1.1 Learning shape Fréchet means
(a) Liver Riemannian regularization.
@@
(c) Rect. Riemannian regularization.
(d) Rect. non-Riemannian regularization.
@@
Figure 5: Template (atlas) building. The learned templates of two different models trained on the rectangles (Rect.) dataset and the liver dataset. We use the model learned using the Riemannian LDDMM regularization (a and c) and the model learned using the non-Riemannian pointwise loss (b and d). We use $\eta=0.05$ and $\eta=50$ in Equation (15) for the rectangles and liver dataset, respectively.
@@
For training the models with Riemannian LDDMM regularization, we need to pick a value for $\eta$. We choose $\eta=0.05$ for the rectangles dataset as we expect rigid body motions from the template to the training shapes. For the liver data, the only prior that we have is that the liver should not change too quickly. Hence, we pick $\eta=50$ such that $\left\lVert\cdot\right\rVert_{V}\approx\eta\left\lVert\cdot\right\rVert_{L_{2}(\Omega)}$.
After training, we obtain the template shapes in Figure 5. While the templates of the liver dataset look almost identical, the templates for the rectangles dataset differ. In particular, both liver templates resemble livers, whereas the template learned on the rectangles data only resembles the data for the Riemannian model. The spherical template in Figure 5d is not desired as it lies outside the data distribution, as depicted in the third category in Figure 2.
@@
The difference between the liver and rectangle datasets stems from the difference in the prior. For the liver dataset, there is no clear deformation prior besides that the liver should not change too quickly. The non-Riemannian regularization induces this prior by construction while the Riemannian regularization enforces this by our choice of $\eta=50$. More precisely, as $\left\lVert\cdot\right\rVert_{V}\approx\eta\left\lVert\cdot\right\rVert_{L_{2}(\Omega)}$ and as the pointwise loss can be interpreted as a non-Riemannian version of the LDDMM regularization with $\left\lVert\cdot\right\rVert_{V}=\left\lVert\cdot\right\rVert_{L_{2}(\Omega)}$ (see Appendix D.2), both regularizations induce the same prior on the deformation, yielding similar templates in Figures 5a and 5b.
@@
While the liver shapes should only change gradually, the rectangles should also move rigidly. Consequently, we let the Killing energy play a more prominent role in $\left\lVert\cdot\right\rVert_{V}$ in the Riemannian LDDMM regularization. This choice yields almost rigid deformations from the template to the training shapes, which results in the square template in Figure 5c. As the pointwise loss does not induce an isometry prior and is a non-Riemannian version of the LDDMM regularization with $\left\lVert\cdot\right\rVert_{V}=\left\lVert\cdot\right\rVert_{L_{2}(\Omega)}$, we obtain the spherical template in Figure 5d.
@@
In summary, while on the liver dataset, the templates look similar, the Riemannian regularization is required for the rectangles dataset to learn a template that fits the data distribution. Hence, the Riemannian LDDMM regularization allows for more flexibility in learning a template shape that resembles the data. Moreover, on data with a clear Riemannian prior that not only promotes small $\left\lVert v\right\rVert_{L^{2}(\Omega)}$, such as a rigidity prior, the Riemannian LDDMM regularization is preferred.
@@ 5.1.2 Learning geodesic shape deformations
Figure 6: Template deformations. The deformation of the template into a reconstructed rectangle using the model learned with the Riemannian LDDMM regularization (top) and with the model learned using the non-Riemannian pointwise loss (bottom). The colors represent the matching of the points to their template.
@@
Figure 7: Template deformations. The deformation of the template into a reconstructed liver using the model learned with the Riemannian LDDMM regularization (top) and with the model learned using the non-Riemannian pointwise loss (bottom). The colors represent the matching of the points to their template.
The previous section shows a regularizer’s effect on the learned Fréchet mean. To finalize the discussion on Figure 2, we assess the variance calculation by qualitatively assessing whether the template deformations are geodesics.
In Figure 6 the template is transformed into a reconstructed training shape for each model trained on the rectangles dataset. We notice that the non-Riemannian model quickly transitions from the spherical template to the target shape and then has a tendency to stay in a nearly identical configuration. In contrast, the model with the Riemannian LDDMM regularization smoothly rotates and scales the rectangular template shape to the final reconstructed shape. Hence, it fits the rigid body motion prior induced by the Killing energy.
@@
Regarding the liver dataset, Figure 7 shows that the deformation with the Riemannian LDDMM regularization is again smoother in time than the deformation with the non-Riemannian pointwise loss. Figure 8 reinforces this as it shows that the vector field of the non-Riemannian model at time $t=0.0$ is everywhere much larger than the vector field of the Riemannian model. Moreover, at $t=0.2$, the vector field of the Riemannian model points in approximately the same direction as at $t=0.0$, while this does not apply to the non-Riemannian model. Hence, the vector field of the Riemannian model is smoother in time.
@@
To explain the preceding observations, we note that, by construction, the non-Riemannian pointwise loss encourages a rapid transition to the target shape followed by a tendency to stay in a nearly identical configuration. Consequently, the velocity vector fields at later times $t$ fine-tune the obtained reconstructions. Hence, the velocity vector field at $t=0.2$ does not necessarily point in approximately the same direction as the velocity vector field at $t=0.0$, as shown in Figure 8. In contrast, the Riemannian LDDMM regularization penalizes rapid transitions and prefers to gradually get closer to the target shape in a prior-consistent manner. Hence, the Riemannian regularization allows for a smooth, physically plausible template deformation into the reconstructed shapes.
Figure 8: Velocity vector fields. The velocity vector fields at $t=0$ and $t=0.2$ (zoom-in) of the deformations in Figure 7. The colors represent the magnitude of the vector field. The non-Riemannian model produces an unrealistic magnitude at $t=0$ and an inconsistent vector field at $t=0.2$ as a consequence.
@@
In Table 2, we summarize the discussion regarding Fréchet means and variance. The table shows that the Riemannian LDDMM regularization is more flexible in influencing the template than the non-Riemannian pointwise loss. Moreover, to obtain good variance estimates, the LDDMM regularization is needed to obtain geodesic deformations of the template to the reconstructions. Hence, the model with the Riemannian LDDMM regularization is the only model in category 1 of Figure 2 and, therefore, is the model with the best mean-variance analysis. The Riemannian model allows us to calculate Fréchet means of the data, to calculate physically plausible geodesic deformations between the template shape and another shape, and to approximate the Riemannian distance between the template and a target shape. Hence, we have added Riemannian geometry to the latent space model.
@@
Table 2: PGA performance. Summary of the results of Sections 5.1.1 and 5.1.2 regarding the models with Riemannian LDDMM regularization and non-Riemannian pointwise regularization. The results relate to Fig. 2.
@@
Data-consistent mean Proper computation variance
@@
Table 3: Training and test set evaluation. The model with the Riemannian LDDMM regularization and the model with the non-Riemannian pointwise regularization are evaluated by reconstructing the training and test sets of the rectangles (Rect.) and liver dataset. We calculate the average value and median (between brackets) of the Chamfer Distance (CD) and Earth Mover distance (EM). The CD values are of the order $10^{-4}$. The best (smallest) values are in bold.
@@
(a) Training set evaluation
@@
(b) Test set evaluation
5.1.3 Generalizability of shape encoding
This section evaluates the reconstruction quality of the two different models by reconstructing the test sets of the rectangles and liver dataset. We use the Chamfer Distance (CD) and the Earth Mover Distance (EM) as evaluation metrics. Table 3a shows the reconstruction metrics of the training data. We see that both models perform approximately equally on the training data.
Table 3b shows the reconstruction metrics on the test sets. While the Riemannian and non-Riemannian methods on the liver dataset only penalize large deviations from the template and obtain similar reconstruction performance, the Riemannian regularization uses a rigidity prior that aligns more closely with the data on the rectangles dataset and obtains improved performance over the non-Riemannian model. This difference demonstrates the importance of a data-consistent prior on deformations to induce a meaningful prior on the latent space that enhances the model’s generalization. All in all, on data with a clear Riemannian prior that not only promotes small $\left\lVert v\right\rVert_{L^{2}(\Omega)}$, such as a rigidity prior, the Riemannian LDDMM regularization improves generalization. Hence, physical consistency can enhance generalization performance.
@@ 5.1.4 Robustness to noise
In this section, we investigate how noise affects the reconstruction performance of the two models. We add random Gaussian noise with mean zero and a standard deviation $\delta$ of $0.01$ or $0.02$ to the vertices of the meshes in the test set. Subsequently, we reconstruct these noisy meshes and compare the reconstructions to the noiseless ground truth meshes. Figure 9 provides some of these reconstructions. In general, both models are effective in denoising the illustrated input meshes. On the rectangles, the Riemannian regularization produces good, noise-resistant reconstructions. The model with the non-Riemannian regularization generates worse reconstructions and is affected by the noise, as illustrated by the reconstructions, which gradually lose their rectangular shape when the noise increases. On the liver dataset, both models produce satisfactory denoised reconstructions. However, while the model with Riemannian regularization yields reconstructions that are approximately independent of the noise level, the model with non-Riemannian regularization results in visibly different reconstructions when the noise level varies.
Figure 9: Robust reconstruction. The reconstruction of two test shapes with added zero-mean Gaussian noise for different standard deviations $\delta$. The red and blue colors represent the matching of the points to the learned template. The Riemannian regularization stabilizes the reconstruction procedure against noise.
To assess whether the observations concerning Figure 9 extend to other liver and rectangle shapes, Table 4 presents the average and median reconstruction errors between the reconstructed shapes and the ground truth noiseless shapes. Comparing the values to Table 3b, we notice that on the rectangles dataset, the model with Riemannian LDDMM regularization is much less sensitive to the noise than the model with the non-Riemannian regularization. On the liver dataset, the median values are relatively stable for both models. Moreover, the average reconstruction errors of both models are not influenced much by a noise level of $0.01$. However, when increasing the standard deviation to $0.02$, the non-Riemannian model attains a much bigger average reconstruction error. In contrast, the Riemannian model attains a similar average error to the scenario with a $0.01$ noise level.
Table 4: Evaluating robustness. The models with the non-Riemannian pointwise and the Riemannian LDDMM regularization are compared by reconstructing noisy test sets ($\delta=0.01$ and $\delta=0.02$) of the rectangles (Rect.) and liver dataset. We calculate the average value and median (between brackets) of the Chamfer Distance (CD) and Earth Mover distance (EM). The values are obtained by comparing the noiseless shapes to the reconstructions of their noisy versions. The CD values are of order $10^{-4}$. The two smallest values are in bold, showing the increased robustness of the Riemannian approach.
@@
The latter observation is related to Figure 9 where the liver reconstructions obtained via the non-Riemannian model depend on the noise level. As discussed in Section 5.1.2, the non-Riemannian regularization encourages a rapid transition to the target shape followed by smaller deformations for fine-tuning. Such large displacements can cause sensitivity to small changes and might yield worse or different reconstructions. Disallowing such large displacements via the Riemannian velocity field regularization stabilizes the problem.
In summary, the physical consistency obtained by the Riemannian regularization improves robustness. The improved robustness is largest on data with a clear Riemannian prior that not only promotes small $\left\lVert v\right\rVert_{L^{2}(\Omega)}$, such as the rigidity prior of the rectangles dataset.